OpenStudy (anonymous):

Where do the graphs of the following equations intersect: x-y=3 x=-2y-6 please someone who is actually willing to help me i struggle on this

OpenStudy (anonymous):

ok. first you need to have one of the equations in terms of x or y. Lets take x-y=3. solve for x ineedyouubiebs

OpenStudy (anonymous):

so it doesnt matter which one i start with?

OpenStudy (anonymous):

no, it does not matter. but x-y=3 is easier to work with.

OpenStudy (anonymous):

Solve for y, x-y=3 -x -y=-x+3 times -1 y=x-3 x=-2y-6 -6 x-6=-2y divde buy -2 -1/2x+3=y or y=-1/2x+3

OpenStudy (anonymous):

would it be (-2y-6)-y=3 ?? please correct me if im wrong

OpenStudy (anonymous):

@ineedyouubiebs , that would be correct... ^^^

OpenStudy (anonymous):

omgggg yesss!!

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

@dpaInc and how would i add like terms to that...i mean would i have to distribute?

OpenStudy (anonymous):

technically you are distributing a "1" in 1(-2y-6)-y=3 but since it is only a positive 1, just get rid of the parenthesis to get: -2y - 6 - y = 3 now combine like terms and solve....

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

correct... now solve for x...

OpenStudy (anonymous):

and how would i do that?

OpenStudy (anonymous):

you have x = -2y - 6 from the beginning....

OpenStudy (anonymous):

just plug in the value that you found for y: x = -2(-3) - 6

OpenStudy (anonymous):

x=6-6

OpenStudy (anonymous):

simplify that....

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

you're basically done... and whenever you need to see if your answer is correct or not, plug in the values of both x and y into BOTH equations. if they both turn out to be true statements, your answer is correct.

OpenStudy (anonymous):

okay thank you soo much=D can you help me in another problem?

OpenStudy (anonymous):

sure.... close this question and post a new one.... :)